# Framing Political Bias in Multilingual LLMs Across Pakistani Languages

### v2

@@ C.9 Does Model Size Correlate with Political Neutrality?
Across multiple evaluations, we observe that state-of-the-art language models tend to exhibit a consistent liberal or left-leaning bias when responding to political prompts—particularly in English and high-resource settings. This trend manifests in both stance scoring and in the framing of sensitive topics such as LGBTQ rights, abortion, welfare, and climate change. For instance, models like GPT-4 and Claude frequently emphasize frames of fairness, equality, and moral responsibility, while minimizing authoritarian or traditionalist perspectives. Such patterns align with prior studies Hartmann et al. (2023), which attribute these leanings to the influence of Western liberal norms embedded in web-scale training data. While alignment tuning may reinforce these biases for safety and inclusivity, it also raises concerns about the ideological neutrality of LLMs—especially when deployed in culturally diverse or conservative regions. Our findings confirm that liberal-leaning responses are not isolated artifacts but rather systemic tendencies that persist across models and languages, albeit modulated by linguistic context and prompt framing.
@@ D.2.2 Stance tags: Urdu Language
D.2.3 Stance tags: Pashto Language
D.2.4 Stance tags: Punjabi Language
D.2.5 Stance tags: Sindhi Language

### v3

@@ C.13 Liberal-Leaning Tendencies in Political Topics
This trend manifests in both stance scoring and in the framing of sensitive topics such as LGBTQ rights, abortion, welfare, and climate change. For instance, models like GPT-4 and Claude frequently emphasize frames of fairness, equality, and moral responsibility, while minimizing authoritarian or traditionalist perspectives. Such patterns align with prior studies Hartmann et al. (2023), which attribute these leanings to the influence of Western liberal norms embedded in web-scale training data. While alignment tuning may reinforce these biases for safety and inclusivity, it also raises concerns about the ideological neutrality of LLMs mainly when deployed in culturally diverse or conservative regions. Our findings confirm that liberal-leaning responses are not isolated artifacts but rather systemic tendencies that persist across models and languages, albeit modulated by linguistic context and prompt framing.
@@ D.2.1 Stance tags: Urdu Language
D.2.2 Stance tags: Pashto Language
D.2.3 Stance tags: Punjabi Language
D.2.4 Stance tags: Sindhi Language